he diameter of a ball bearing was measured by 12 inspectors, each using two different kinds of calipers. The results were Caliper 1 1.238 1.238 1.239 1.24 1·24 1.238 1.24 1 .24 1.238 1.241 1.241 1.238 Caliper 2 1.227 1.228 1.227 1.229 1.23 1.231 1.227 1.228 1.228 1·23 1.231 1.232 Inspector 2 4 7 12 Is there a significant difference between the mean diameter measurement of caliper 1 and caliper 2? Conduct the appropriate hypothesis test using α-0.05.
Answer #1

Hypothesis:

Ho : μ1 = μ2 Or There is No significant difference between the mean diameter measurement of caliper 1 and caliper 2

Ha: \mu _{1} \neq \mu _{2} Or There is significant difference between the mean diameter measurement of caliper 1 and caliper 2

Calculation:

For caliper 1

Mean = X1-1.24 , Standard deviation (S1) = 0.0012, n1 = 12

For caliper 2

Mean = X2 = 1.23 , Standard deviation (S2) = 0.0018, n2 = 12

Test statistic:

1.24 1.23 16.61 S12S22 0.001220.00182 12 12

Degrees of Freedom:

0.00122 0 0.00182 12 0.00122 12-1 7 12 df 20 0.00182 n1-1 12-1

Critical value:

t_{\alpha/2 ,df} = t_{0.05/2 ,20} = 2.0860 ...................By using t table

Conclusion:

Here, Test statistic > Critical value, i.e. 16.61 > 2.0860, That is Reject Ho at 5% level of significance.

Therefore,

There is significant difference between the mean diameter measurement of caliper 1 and caliper 2